OpenStudy (anonymous):

how do i find the missing exponent to (6)(6^3)=6^5

OpenStudy (anonymous):

What do you mean missing exponent? If you're multiplying a number by itself raised to some power, you just add one to the power. So \(6\times6^3 = 6^4\)

OpenStudy (anonymous):

(6^)(6^3)=6^5

OpenStudy (anonymous):

Ok, well, the idea still holds. How many times do you have to multiply \(6^3\) by 6 to get \(6^5\)?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Right. And an exponent is just multiplying a number by itself that many times. So you have your answer.

OpenStudy (anonymous):

thanks